# Uncertainty relations for overcomplete measurements from generalized equiangular tight frames

### v1

@@ Abstract
The current study aims to examine uncertainty relations for overcomplete measurements from generalized equiangular tight frames. Informationally overcomplete measurements are a valuable tool in quantum information processing, including tomography and state estimation. The maximal sets of mutually unbiased bases are the most common case of informationally overcomplete measurements. The existence of $d+1$ of mutually unbiased bases is proved for $d$ being a prime power. More general classes of informationally overcomplete measurements have been proposed for various purposes. Such measurements are typically characterized by some inner structure maintaining the required properties. It leads to restrictions imposed on generated probabilities. To apply the measurements of interest, these restrictions should be converted into information-theoretic terms. To describe the amount of uncertainty quantitatively, we use the Tsallis and Rényi entropies as well as probabilities of separate outcomes. The obtained results are based on estimation of the index of coincidence. The derived relations are briefly exemplified.
overcomplete measurements, index of coincidence, Tsallis entropy, Rényi entropy, conical designs
@@ I Introduction
In fact, the problem of building desired structures in Hilbert spaces is often hard to resolve [8, 16]. It is known that $d+1$ of mutually unbiased bases in dimension $d$ exist for $d$ being a prime power. But the maximal number of mutually unbiased bases remains unknown even for $d=6$ [17]. There are reasons to believe that SIC-POVMs exist for all $d$, and this conjecture is due to Zauner [18]. Here, there are ways to weak some of the imposed restrictions. Dealing with rank-one measurement operators, we may reduce the number of outcomes. In this regard, equiangular tight frames are shown to deserve more attention than they have obtained [19]. Equiangular measurements are a special case of more general concepts such as equioverlapping measurements [20, 21] and the so-called $(1,M)$-POVMs [22]. Another way is to allow measurement operators of arbitrary rank. General symmetric informationally complete measurements were built [23, 24] and applied [25]. Mutually unbiased measurements instead of bases were proposed in [26]. Recently, generalized symmetric measurements [27] and generalized equiangular ones [28] have been studied.
@@
The Heisenberg uncertainty principle [29] is widely recognized as a fundamental concept that the observer perturbs what he observes. Since the first formal derivations of Kennard [30] and Robertson [31] appeared, many approaches and scenarios were addressed [32]. For finite-dimensional systems, entropic uncertainty relations [33, 34] give a useful alternative to Robertson’s formulation. The well-known Maassen–Uffink uncertainty relation [34] was originally conjectured by Kraus [35]. This relation becomes a tool in building feasible schemes to detect of nonclassical correlations such as entanglement [36, 37] and steerability [38, 39, 40]. Entropic uncertainty relations have also emerged in the security analysis of quantum cryptographic protocols [41, 42, 43]. Entropic formulation gives a natural way to address uncertainties in the presence of quantum memory [44]. For more results on entropic uncertainty relations, see the reviews [45, 46, 47] and references therein. For measurements with a special inner structure, the Maassen–Uffink uncertainty relation can be improved [48, 49, 50, 51].
@@
It is important to examine properties of specially built quantum measurements as completely as possible, even if the final opinion on their use is now unformed. The aim of this study is to derive uncertainty relations for overcomplete measurements from generalized equiangular tight frames. It immediately concerns uncertainty relations for generalized symmetric measurements due to the correspondence mentioned in [28]. The new relations are expressed in terms of the Renyi [52] and Tsallis entropies [53] as well as sums of probabilities. Some of the used inequalities are based on considering information diagrams as described in [54, 55]. The paper is organized as follows. Section II introduces the preliminary material concerning generalized equiangular measurements and generalized symmetric ones. Some properties of the Tsallis and Rényi entropies are recalled as well. New uncertainty relations for the considered measurements are presented in Sec. III. A utility of the obtained inequalities is briefly exemplified. Section IV concludes the paper with a summary of the results.
@@ II Preliminaries
This section aims to recall the required material concerning informationally overcomplete measurements derived from generalized equiangular tight frames. It also considers entropies that will be used to characterize uncertainties in quantum measurements of interest.
@@ II.1 Generalized equiangular measurements and conical $2$-designs
Basic reasons to put informationally overcomplete measurements are discussed in [28]. We recall only very necessary facts about them. Let ${\mathcal{H}}_{d}$ be $d$-dimensional Hilbert space. By ${\mathcal{L}}({\mathcal{H}}_{d})$ and ${\mathcal{L}}_{+}({\mathcal{H}}_{d})$, we respectively mean the space of linear operators on ${\mathcal{H}}_{d}$ and the set of positive semidefinite ones. The state of a quantum system is described by density matrix $\boldsymbol{\rho}\in{\mathcal{L}}_{+}({\mathcal{H}}_{d})$ normalized as $\mathrm{tr}(\boldsymbol{\rho})=1$. In general, any quantum measurement is represented as a set =${\mathcal{E}}=\bigl{\{}{\mathsf{E}}_{j}\bigr{\}}_{j=1}^{N}$ with ${\mathsf{E}}_{j}\in{\mathcal{L}}_{+}({\mathcal{H}}_{d})$ such that the completeness relation holds,
@@
For the pre-measurement state $\boldsymbol{\rho}$, the probability of $j$-th outcome is equal to $\mathrm{tr}({\mathsf{E}}_{j}\boldsymbol{\rho})$. In principle, the number of different outcomes can exceed the dimensionality. This property is indispensable in many questions of quantum information science.
@@
$\mathrm{tr}({\mathsf{P}}_{\mu,i}{\mathsf{P}}_{\nu,j})$ $=f\,\mathrm{tr}({\mathsf{P}}_{\mu,j})\,\mathrm{tr}({\mathsf{P}}_{\nu,j})\qquad(\mu\neq\nu)\,.$ (7)
@@
The admissible range for $b_{\mu}$ is obtained as [28]
$\frac{1}{d}<b_{\mu}\leq\frac{\min\{d,N_{\mu}\}}{d}\ .$ (9)
A single equiangular measurement is a particular case for $M=\gamma=b=1$. It should be noted here that our notation slightly differs from the notation of [28]. The generalized equiangular measurement can also be interpreted as a set of $M$ POVMs so that $\mu$-th POVM has elements $\gamma_{\mu}^{-1}\,{\mathsf{P}}_{\mu,j}$. Due to this correspondence, a generalized equiangular measurement relates to some generalized symmetric measurement [28].
The generalized symmetric measurements were introduced as collections of POVMs that are not equinumerous [27]. It is important here that elements of generalized equiangular measurements can be constructed from elements of generalized symmetric measurements. One uses [27] a set of $M$ POVMs ${\mathcal{E}}_{\mu}=\bigl{\{}{\mathsf{E}}_{\mu,j}\bigr{\}}_{j=1}^{N_{\mu}}$ satisfying additional symmetry conditions:
@@
There are four construction methods with the use of linear combinations of Hermitian orthonormal operator bases [27]. We refrain from presenting the details here.
The following observation has been made in [28]. Suppose that positive numbers $\gamma_{\mu}$ obey (3) and $M$ POVMs ${\mathcal{E}}_{\mu}=\bigl{\{}{\mathsf{E}}_{\mu,j}\bigr{\}}_{j=1}^{N_{\mu}}$ satisfy (10)–(13). Then operators of the form
@@
compose a generalized equiangular measurement. The correspondence between the coefficients reads as
@@
The obtained POVM is informationally complete, if it is constructed from informationally complete set of elements ${\mathsf{E}}_{\mu,j}$ and $\gamma_{\mu}>0$ for all $\mu=1,\ldots,M$ [28]. Thus, the generalized symmetric measurements are sufficient to build a generalized equiangular measurement. Despite of simplicity of the relation (15), this construction has some nice features. For example, it is possible to manipulate the coefficients in such a way that both the traces $\mathrm{tr}({\mathsf{P}}_{\mu,j}^{2})$ and $\mathrm{tr}({\mathsf{P}}_{\mu,i}{\mathsf{P}}_{\mu,j})$ become $\mu$-independent [28].
Projective designs have found use in various questions of quantum information theory. It seems that the special case of $2$-designs is provided with a large number of examples. Let $\bigl{\{}|\phi_{j}\rangle\bigr{\}}_{j=1}^{N}$ be a set of $N$ unit vectors of ${\mathcal{H}}_{d}$. There are several equivalent definitions, when these vectors form a projective $2$-design [18, 56]. One of them reads as
$\frac{1}{N}\,\sum_{j=1}^{N}|\phi_{j}\rangle\langle\phi_{j}|\otimes|\phi_{j}\rangle\langle\phi_{j}|=\frac{2}{d^{2}+d}\>{\mathsf{\Pi}}_{\mathrm{sym}}\,,$ (17)
where ${\mathsf{\Pi}}_{\mathrm{sym}}$ is the orthogonal projection on the symmetric subspace of ${\mathcal{H}}_{d}\otimes{\mathcal{H}}_{d}$. Conical designs are generalizations of the above concept to positive operators [57, 58]. We introduce the unitary swap operator $\tilde{\mathsf{W}}$ on ${\mathcal{H}}_{d}\otimes{\mathcal{H}}_{d}$ which takes $|\psi\rangle\otimes|\varphi\rangle$ to $|\varphi\rangle\otimes|\psi\rangle$. The set $\bigl{\{}{\mathsf{A}}_{j}\bigr{\}}_{j=1}^{N}$ with ${\mathsf{A}}_{j}\in{\mathcal{L}}_{+}({\mathcal{H}}_{d})$ is a conical $2$-design, if and only if
@@
where $\kappa_{+}\geq\kappa_{-}>0$. It can be noticed that $2\kappa_{\pm}=\kappa_{\mathrm{s}}\pm\kappa_{\mathrm{a}}$. Taking the partial trace in (18) yields
[…]
To each conical $2$-design, we can assign POVM =${\mathcal{E}}=\bigl{\{}{\mathsf{E}}_{j}\bigr{\}}_{j=1}^{N}$ with elements
${\mathsf{E}}_{j}=\frac{\mathrm{tr}({\mathsf{A}}_{j})}{\kappa_{+}d+\kappa_{-}}\,{\mathsf{A}}_{j}\,.$ (20)
The case with constant $\mathrm{tr}({\mathsf{A}}_{j})$ is very close to a projective $2$-design. On the other hand, a possibility to manipulate the coefficients generally excludes this restriction.
The author of [28] showed that some of the generalized equiangular measurements are conical $2$-designs. Let us take the coefficients such that
@@
There exist conical $2$-designs even though they are constructed from the generalized symmetric measurements that are not conical $2$-designs [28]. This property could be useful in producing new conical $2$-designs. Unitary designs are now considered as a powerful tool in quantum information science [59, 60]. Thus, generalized equiangular measurements deserve to be studied in application to questions of quantum information science. It turned out that generated probabilities satisfy certain conditions that lead to various restrictions including uncertainty relations. To express the amount of uncertainties, the Tsallis and Rényi entropies will be used in this paper.
@@ II.2 Generalized entropies and the index of coincidence
The index of coincidence is used in various questions of information theory [61]. Let $P=(p_{1},\ldots,p_{N})$ be a probability distribution with the index running over the set $\{1,\ldots,N\}$. The corresponding index of coincidence reads as
@@
Indicies with larger degrees of probabilities have also found use [61]. Applications of such indicies to derive uncertainty relations were given in [62, 63, 64]. For $\alpha>0$, the Tsallis $\alpha$-entropy is expressed as [53]
@@
It was demonstrated for information diagrams with the Shannon entropy that the theoretically best lower-bounding curve is not smooth [61]. Expectedly, the same feature takes place for the Tsallis $\alpha$-entropy. It can be shown that a smoothness is violated in the points with abscissas $X=k^{-1}$. In more detail, this question is explained in [54, 55]. Actually, the inequality (29) is improved by replacing the graph of smooth function $X\mapsto\ln_{\alpha}\bigl{(}X^{-1}\bigr{)}$ with the polygonal line connecting the points with abscissas $X=k^{-1}$ and ordinates $\ln_{\alpha}(k)$ for integer $k\geq 1$. New graph shows the piecewise linear function $X\mapsto{L}_{\alpha}(X)$ expressed as [54]
@@
This inequality is not tight, except for points of the form $\bigl{(}k^{-1},\ln_{\alpha}(k)\bigr{)}$. For $\alpha=1$, the inequality (33) reduces to one of the main results of [61]. We will apply (33) to derive uncertainty relations for informationally overcomplete measurements and conical $2$-designs.
@@
One of the results of [55] allows one to estimate sums of two probabilities from above at the given index of coincidence. Let the index of coincidence of the distribution $P=(p_{1},\ldots,p_{N})$ obey $N^{-1}\leq{I}(P)\leq 2^{-1}$. It then holds that
@@
The inequality (36) is formally correct for all acceptable values of $I(P)$, but a nontrivial bound is obtained only for $I(P)<2^{-1}$. As a rule, the number of outcomes in overcomplete measurements is sufficiently large. Hence, the last inequality is typically satisfied. For example, this holds for a complete set of $d+1$ MUBs in dimension $d\geq 3$ [55].

### v3

# Uncertainty relations for quantum measurements from generalized equiangular tight frames

@@ Abstract
The current study aims to examine uncertainty relations for measurements from generalized equiangular tight frames. Informationally overcomplete measurements are a valuable tool in quantum information processing, including tomography and state estimation. The maximal sets of mutually unbiased bases are the most common case of such measurements. The existence of $d+1$ mutually unbiased bases is proved for $d$ being a prime power. More general classes of informationally overcomplete measurements have been proposed for various purposes. Measurements of interest are typically characterized by some inner structure maintaining the required properties. It leads to restrictions imposed on generated probabilities. To apply the considered measurements, these restrictions should be converted into information-theoretic terms. It is interesting that certain restrictions hold irrespectively to overcompleteness. To describe the amount of uncertainty quantitatively, we use the Tsallis and Rényi entropies as well as probabilities of separate outcomes. The obtained results are based on estimation of the index of coincidence. The derived relations are briefly exemplified.
symmetric measurements, Tsallis entropy, Rényi entropy, conical designs
@@ I Introduction
In fact, the problem of building desired structures in Hilbert spaces is often hard to resolve [8, 16]. It is known that $d+1$ mutually unbiased bases in dimension $d$ exist for $d$ being a prime power. But the maximal number of mutually unbiased bases remains unknown even for $d=6$ [17]. There are reasons to believe that SIC-POVMs exist for all $d$, and this conjecture is due to Zauner [18]. Here, there are ways to weak some of the imposed restrictions. Dealing with rank-one measurement operators, we may reduce the number of outcomes. In this regard, equiangular tight frames are shown to deserve more attention than they have obtained [19]. Equiangular measurements are a special case of more general concepts such as equioverlapping measurements [20, 21] and the so-called $(N,M)$-POVMs [22, 23]. Another way is to allow measurement operators of arbitrary rank. General symmetric informationally complete measurements were built [24, 25] and applied [26]. Mutually unbiased measurements instead of bases were proposed in [27]. Recently, generalized symmetric measurements [28] and generalized equiangular ones [29] have been studied.
@@
The Heisenberg uncertainty principle [30] is widely recognized as a fundamental concept that the observer perturbs what he observes. Since the first formal derivations of Kennard [31] and Robertson [32] appeared, many approaches and scenarios were addressed [33]. For finite-dimensional systems, entropic uncertainty relations [34, 35] give a useful alternative to Robertson’s formulation. The well-known Maassen–Uffink uncertainty relation [35] was originally conjectured by Kraus [36]. This relation becomes a tool in building feasible schemes to detect nonclassical correlations such as entanglement [37, 38] and steerability [39, 40, 41]. Entropic uncertainty relations have also emerged in the security analysis of quantum cryptographic protocols [42, 43, 44]. Entropic formulation gives a natural way to address uncertainties in the presence of quantum memory [45]. For more results on entropic uncertainty relations, see the reviews [46, 47, 48] and references therein. For measurements with a special inner structure, the Maassen–Uffink uncertainty relation can be improved [49, 50, 51, 52].
@@
It is important to examine properties of specially built quantum measurements as completely as possible, even if the final opinion on their use is now unformed. The aim of this study is to derive uncertainty relations for quantum measurements from generalized equiangular tight frames. In contrast to the treatment of [29], these measurements are not assumed to be overcomplete. The new relations are expressed in terms of the Rényi [53] and Tsallis entropies [54] as well as sums of probabilities. Some of the used inequalities are based on considering information diagrams as described in [55, 56]. The paper is organized as follows. Section II introduces the preliminary material concerning generalized equiangular measurements and generalized symmetric ones. Some properties of the Tsallis and Rényi entropies are recalled as well. New uncertainty relations for the considered measurements are presented in Sec. III. A utility of the obtained inequalities is briefly exemplified. Section IV concludes the paper with a summary of the results. Appendix A deals with a statement concerning indices of coincidence.
@@ II Preliminaries
This section aims to recall the required material concerning quantum measurements derived from generalized equiangular tight frames. It also considers entropies that will be used to characterize uncertainties in quantum measurements of interest.
@@ II.1 Generalized equiangular measurements and conical $2$-designs
Basic reasons to put informationally overcomplete measurements are discussed in [29]. We recall only very necessary facts about them. Let ${\mathcal{H}}_{d}$ be $d$-dimensional Hilbert space. By ${\mathcal{L}}({\mathcal{H}}_{d})$ and ${\mathcal{L}}_{+}({\mathcal{H}}_{d})$, we respectively mean the space of linear operators on ${\mathcal{H}}_{d}$ and the set of positive semidefinite ones. The state of a quantum system is described by density matrix $\boldsymbol{\rho}\in{\mathcal{L}}_{+}({\mathcal{H}}_{d})$ normalized as $\mathrm{tr}(\boldsymbol{\rho})=1$. In general, any quantum measurement is represented as a set =${\mathcal{E}}=\bigl{\{}{\mathsf{E}}_{j}\bigr{\}}_{j=1}^{N}$ with ${\mathsf{E}}_{j}\in{\mathcal{L}}_{+}({\mathcal{H}}_{d})$ such that the completeness relation holds [57]
@@
For the pre-measurement state $\boldsymbol{\rho}$, the probability of $j$-th outcome is equal to $\mathrm{tr}({\mathsf{E}}_{j}\boldsymbol{\rho})$. It is very important that the number of different outcomes can exceed the dimensionality.
@@
$\mathrm{tr}({\mathsf{P}}_{\mu,i}{\mathsf{P}}_{\nu,j})$ $=f\,\mathrm{tr}({\mathsf{P}}_{\mu,i})\,\mathrm{tr}({\mathsf{P}}_{\nu,j})\qquad(\mu\neq\nu)\,.$ (7)
@@
A single equiangular measurement with rank-one elements is a particular case for $M=\gamma=b=1$. The generalized equiangular measurement can also be interpreted as a set of $M$ POVMs so that $\mu$-th POVM has elements $\gamma_{\mu}^{-1}\,{\mathsf{P}}_{\mu,j}$. Due to this correspondence, a generalized equiangular measurement relates to some generalized symmetric measurement [29].
The generalized symmetric measurements were introduced as collections of POVMs that are not equinumerous [28]. A generalized equiangular measurement can be constructed from elements of generalized symmetric measurements. One uses [28] a set of $M$ POVMs ${\mathcal{E}}_{\mu}=\bigl{\{}{\mathsf{E}}_{\mu,j}\bigr{\}}_{j=1}^{N_{\mu}}$ satisfying additional symmetry conditions:
@@
Let positive numbers $\gamma_{\mu}$ obey (3) and $M$ POVMs ${\mathcal{E}}_{\mu}=\bigl{\{}{\mathsf{E}}_{\mu,j}\bigr{\}}_{j=1}^{N_{\mu}}$ satisfy (9)–(12). Then operators of the form
@@
compose a generalized equiangular measurement [29]. The correspondence between the coefficients reads as
@@
The obtained POVM is informationally complete, if it is constructed from informationally complete set of elements ${\mathsf{E}}_{\mu,j}$ and $\gamma_{\mu}>0$ for all $\mu=1,\ldots,M$ [29]. Despite of simplicity of the relation (14), this construction has some nice features. For example, one can manipulate the coefficients so that both the traces $\mathrm{tr}({\mathsf{P}}_{\mu,j}^{2})$ and $\mathrm{tr}({\mathsf{P}}_{\mu,i}{\mathsf{P}}_{\mu,j})$ become $\mu$-independent [29].
Under additional conditions on the coefficients, a generalized equiangular measurement is a conical $2$-design [29]. Projective and conical designs have found use in various questions of quantum information theory. Let $\bigl{\{}|\phi_{j}\rangle\bigr{\}}_{j=1}^{K}$ be a set of $K$ unit vectors of ${\mathcal{H}}_{d}$. There are several equivalent definitions, when these vectors form a projective $2$-design [18, 58]. One of them reads as
$\frac{1}{K}\,\sum_{j=1}^{K}|\phi_{j}\rangle\langle\phi_{j}|\otimes|\phi_{j}\rangle\langle\phi_{j}|=\frac{2}{d^{2}+d}\>{\mathsf{\Pi}}_{\mathrm{sym}}\,,$ (16)
where ${\mathsf{\Pi}}_{\mathrm{sym}}$ is the orthogonal projection on the symmetric subspace of ${\mathcal{H}}_{d}\otimes{\mathcal{H}}_{d}$. Conical designs are generalizations of the above concept to positive operators [59, 60]. We introduce the unitary swap operator $\tilde{\mathsf{W}}$ on ${\mathcal{H}}_{d}\otimes{\mathcal{H}}_{d}$ which takes $|\psi\rangle\otimes|\varphi\rangle$ to $|\varphi\rangle\otimes|\psi\rangle$. The set $\bigl{\{}{\mathsf{A}}_{j}\bigr{\}}_{j=1}^{K}$ with ${\mathsf{A}}_{j}\in{\mathcal{L}}_{+}({\mathcal{H}}_{d})$ is a conical $2$-design, if and only if
@@
where $\kappa_{+}\geq\kappa_{-}>0$. It can be noticed that $2\kappa_{\pm}=\kappa_{\mathrm{s}}\pm\kappa_{\mathrm{a}}$. To obtain a conical $2$-design from equiangular lines, the coefficients should satisfy [29]
@@
There exist conical $2$-designs even though they are constructed from the generalized symmetric measurements that are not conical $2$-designs [29]. This property is useful in producing new conical $2$-designs. Unitary designs are also considered as a powerful tool in quantum information science [61, 62]. It turned out that generated probabilities satisfy certain conditions that lead to various restrictions including uncertainty relations. To express the amount of uncertainties, the Tsallis and Rényi entropies will be used in this paper.
@@ II.2 Generalized entropies and the index of coincidence
The index of coincidence is used in various questions of information theory [63]. Let $P=(p_{1},\ldots,p_{N})$ be a probability distribution with the index running over the set $\{1,\ldots,N\}$. The index of coincidence reads as
@@
Indices with larger degrees of probabilities have also found use [63]. Applications of such indices to derive uncertainty relations were given in [64, 65, 66]. For $\alpha>0$, the Tsallis $\alpha$-entropy is expressed as [54]
@@
It was demonstrated for information diagrams with the Shannon entropy that the theoretically best lower-bounding curve is not smooth [63]. Naturally, the same feature takes place for the Tsallis $\alpha$-entropy. It can be shown that a smoothness is violated in the points with abscissas $X=k^{-1}$. In more detail, this question is explained in [55, 56]. Actually, the inequality (26) is improved by replacing the graph of smooth function $X\mapsto\ln_{\alpha}\bigl{(}X^{-1}\bigr{)}$ with the polygonal line connecting the points with abscissas $X=k^{-1}$ and ordinates $\ln_{\alpha}(k)$ for integer $k\geq 1$. New graph shows the piecewise linear function $X\mapsto{L}_{\alpha}(X)$ expressed as [55]
@@
This inequality is not tight, except for points of the form $\bigl{(}k^{-1},\ln_{\alpha}(k)\bigr{)}$. For $\alpha=1$, the inequality (30) reduces to one of the main results of [63]. We will apply (30) to derive uncertainty relations for quantum measurements of interest including conical $2$-designs.
@@
One of the results of [56] allows us to estimate sums of two probabilities from above at the given index of coincidence. Let the index of coincidence of the distribution $P=(p_{1},\ldots,p_{N})$ obey $N^{-1}\leq{I}(P)\leq 2^{-1}$. It then holds that
@@
The inequality (33) is formally correct for all acceptable values of $I(P)$, but a nontrivial bound is obtained only for $I(P)<2^{-1}$. This inequality takes place for sufficiently large number of outcomes. For example, it holds for a complete set of $d+1$ MUBs in dimension $d\geq 3$ [56].